OpenStudy (shamil98):

How do i find the derivative of (x^2 -3x)e^x

OpenStudy (shamil98):

\[(x^2 -3x)e^x\]

OpenStudy (amistre64):

use product rule

OpenStudy (anonymous):

Use the Product Rule: Keep in mind that the derivative of (x^2 - 3x) is just 2x-3.

OpenStudy (shamil98):

What would I do with the e^x ?..

OpenStudy (shamil98):

Oh, I'm still learning calc from a book and online sources..

OpenStudy (amistre64):

a(x) e^x a'(x) e^x + a(x) e'^x a'(x) e^x + a(x) e^x e^x(a'+a)

OpenStudy (anonymous):

You have two functions, x^2 - 3x and another function called e^x. To find the derivative of the product of two functions, use the Product Rule, which says that if you have two functions, f(x) and g(x), te derivative of f(x)g(x) is f(x)g'(x) + g(x)f'(x)

OpenStudy (shamil98):

F(x) =(x^2-3x) f' = 2x - 3 g(x) = e^x g' = e^x So the answer would then be.. \[(x^2-3x)e^x + (2x-3)e^x\] right?..

OpenStudy (amistre64):

yep